If an employee deposits Rs. 2,000 at the end of each year into his company’s plan which pays 7% interest compounded quarterly, how much will he have in the account at the end of 5 years?

Answer by Theo(13342)

use the formula f = p * (1 + r) ^ n

interest rate per time period = 7% per year / 4 = 1.75% per quarter / 100 = .0175 per quarter.

2000 is deposited at the end of each year.

future value of 2000 deposited at the end of year 1 = 2000 * (1.0175) ^ 16 = 2639.858702.

future value of 2000 deposited at the end of year 2 = 2000 * (1.0175) ^ 12 = 2462.87863.

future value of 2000 deposited at the end of year 3 = 2000 * (1.0175) ^ 8 = 2297.763566.

future value of 2000 deposited at the end of year 4 = 2000 * (1.0175) ^ 4 = 2143.718063.

future value of 2000 deposited at the end of year 5 = 2000 * (1.0175) ^ 0 = 2000.

total future value at the end of year 5 = 11,544.21896.

this can also be done using a financial calculator to give you future value of payment at the end of each year.

since the interest rate is compounded, you need to find the effective annual interest rate and use that.

the effective annual interest rate would be (1 + .0175) ^ 4 = 1.07185903129 - 1 = .071859031 * 100 = 7.185903129%

using the TI-BA-II financial calculator, i made the following entries.

present value = 0
future value = 0
payment per year = 2000
effective interest rate per year = 7.185903129%
payments are made at the end of each year.

the calculator then computed the future value to be equal to 11,544.21896.

the results from the manual calculation and the financial calculator calculations both agree.

when the calculations are done by formula, the interest rate is used.
the interest rate is the percent interest rate divided by 100.

when the calculator is used, the percent interest rate is used.
the percent interest rate is the interest rate multiplied by 100.